# Formal conjugacy and asymptotic differential algebra

### v3

@@ Abstract
We study conjugacy of formal derivations on fields of generalised power series in characteristic 00 and dimension $1$. Casting the problem of Poincaré resonance in terms of asymptotic differential algebra, we give conditions for conjugacy of parabolic flat $\log$-$\exp$ transseries, flat grid-based transseries, logarithmic transseries, power series with exponents and coefficients in an ordered field, and formal Puiseux series.
@@ Introduction
In this paper, we focus on the conjugacy problem for parabolic series in differential valued fields [34]. These include for instance formal Laurent or Puiseux series, $\log$-$\exp$ transseries or grid-based transseries [24], logarithmic transseries [21], or finitely nested hyperseries [5], and complexifications thereof. This choice is not fortuitous but motivated by the connections between the setting of (ordered) asymptotic differential algebra and that of (ordered) valued groups (see [6, Remark 7.27]). One of the difficulties of solving conjugacy equations for formal series endowed with a composition law $\circ$ and a derivation $\partial$ is that this requires a good understanding of the interaction of the composition law with the valuation. On this path, one is confronted with intricate and computationally heavy problems involving Taylor expansions of arbitrarily high orders, monotonicity of the composition law, and mean value inequalities. It is not the least of hindrances that such properties of $\circ$ and $\partial$ may not have been established for the given algebra of formal series.
We circumvent these issues by leveraging the Lie-type correspondence, given by a formal exponential map $\exp$, between near-identity substitutions $f\mapsto f\circ(x+\delta)$ and contracting derivations $f\mapsto g\partial(f)$ on algebras of formal series. We showed [9] that a fraction of the theory of Lie groups applies to such algebras. We focus on contracting derivations on fields of generalised series, corresponding on the side of vector fields to the suitable generalisation of nilpotent vector fields (see [28]).
@@
[Theorem 2.1] Suppose that $\mathbb{S}$ has regular asymptotic integration. Let $f,g\in\operatorname{Cont}(\partial)\setminus\{0\}$. Then $f$ and $g$ are conjugate in $\operatorname{Cont}(\partial)$ if and only if $v(f-g)>v(g)$ and $v(f-g)>v\left(\left(h/gh^{\prime}\right)^{\prime}\right)$ for all $h\in\mathbb{S}^{\times}$ with $v(h)>0$.
@@
[Theorem 2.2] Suppose that $\mathbb{S}$ has regular asymptotic integration. Let $f,g\in\operatorname{Cont}(\partial)\setminus\{0\}$ such that $v(f-g)>v(f)$ and and $v(f-g)>v\left(\left(h/gh^{\prime}\right)^{\prime}\right)$ for all $h\in\mathbb{S}^{\times}$ with $v(h)>0$. Then the derivations $f\partial$ and $g\partial$ are conjugate over $\operatorname{Aut}(\mathbb{S})$, i.e. there is a $\sigma=\exp(h\partial)\in\operatorname{Aut}(\mathbb{S})$ such that $\sigma\circ(g\partial)\circ\sigma^{\operatorname{inv}}=f\partial$.
In certain cases, the group of automorphisms $\exp(\operatorname{Cont}(\partial))$ is isomorphic to a well-identified group of series under composition. Let $\mathbb{S}$ be the field of transseries whose transmonomials $\mathfrak{m}$ satisfy $v(\mathfrak{m}^{\prime}/\mathfrak{m})+v(x)\geqslant 0$. We identify (Proposition 3.1) the group $\exp(\operatorname{Cont}(\partial))$ for all direct limits of subsystems (Definition 3.1) of the direct system which defines $\mathbb{S}$. Combining this with 1, we obtain:
@@ 1 Groups of contracting derivations
Throughout the paper, we fix a field $C$ of characteristic 00, a non-empty directed set $(D,\leqslant)$ and a directed system $\mathcal{S}=(\Gamma_{d})_{d\in D}$, for the inclusion, of non-trivial ordered Abelian groups. We write $\Gamma$ for the direct limit of $\Gamma_{d}$.
@@
Lastly, we write $1$-$\operatorname{Aut}^{\mathcal{S}}(\mathbb{S})$ for the group, under composition, of regular algebra automorphisms $\sigma$ of $\mathbb{S}$ such that $\sigma(f)-f\prec f$ for all $f\in\mathbb{S}^{\times}$. Since each $C\operatorname{Id}_{\mathbb{S}_{d_{0}}}+\operatorname{Lin}^{+}_{\prec}(\mathbb{S}_{d_{0}})$ has evaluations in the sense of [9, Definition 2.3], and in view of the definition of regularity, so has $C\operatorname{Id}_{\mathbb{S}}+\operatorname{Lin}^{\mathcal{S}}_{\prec}(\mathbb{S})$. So [9, Theorem 2.14] applies and yields:
@@
$\exp:\operatorname{Der}^{\mathcal{S}}_{\prec}(\mathbb{S})$$\longrightarrow$ $1\text{-}\operatorname{Aut}^{\mathcal{S}}(\mathbb{S})$
[…]
Hidden terms in Equation 1 are $\mathbb{Q}$-linear combinations of iterated Lie brackets of lengths > $>\>3$.
@@ 1.1 Contractive hull of a derivation
The inverse of an $f\in\operatorname{Cont}(\partial)$ for $\ast$ is simply $-f$. We also have the following consequences of [6, Lemmas 7.14 and 7.15 and Remark 7.19]. We give the proofs here for completion.
@@
For $f,g\in\operatorname{Cont}(\partial)\backslash\{0\}$, we have $\llbracket f,g\rrbracket\prec f,g$.
@@
In view of Equation 2, Lemmas 1.1 and 1.2 and Remark 1.1, we obtain:
Corollary 1.1
For $f,g\in\operatorname{Cont}(\partial)\backslash\{0\}$, we have $f\ast g\sim f+g$.
@@
If $D=\{\bullet\}$, $\Gamma_{\bullet}=\Gamma=(\mathbb{Z},+,0,<)$, and $\partial=\frac{\mathrm{d}}{\mathrm{d}t}$ is the derivation with respect to $t$ on $C\left(\!\!(t)\!\!\right)=C\left(\!\!(\mathbb{Z})\!\!\right)$, then $\operatorname{Cont}(\partial)=\left\{f\in C\left(\!\!(\mathbb{Z})\!\!\right):f\prec t^{2}\right\}$.
@@ 1.2 Integration and asymptotic integration
$f\prec g\Longleftrightarrow f^{\prime}\prec g^{\prime}\text{{\hskip 30.00005pt}and{\hskip 30.00005pt}$f\prec g\Longrightarrow f^{{\dagger}}\succcurlyeq g^{{\dagger}}.$}$
@@
$\mathord{{}^{\prime}}:\Gamma\setminus\{0\}\longrightarrow\Gamma\>;\>v(g)\mapsto v(g^{\prime})\text{{\hskip 30.00005pt}and{\hskip 30.00005pt}$\mathord{\>{}^{{\dagger}}}:\Gamma\setminus\{0\}\longrightarrow\Gamma\>;\>v(g)\mapsto v(g^{{\dagger}})$},$
@@
A crucial property of $(\mathbb{S},\preccurlyeq,\partial)$ is that [4, Theorem 9.2.1] there is at most one $\gamma\in\Gamma$ such that $\gamma\not\in(\Gamma\setminus\{0\})^{\prime}$. If no such $\gamma$ exists, then $\mathbb{S}$ is closed under asymptotic integration. Indeed, since $\operatorname{Ker}(\partial)=C$, for any $f\in\mathbb{S}\setminus\{0\}$ and any $g$ with $v(g^{\prime})=v(f)$, the element $cg$ is an asymptotic integral of $f$ where $c$ is the leading coefficient of $f/g^{\prime}$. If such an element exist, then we call it the pseudo-gap of $(\mathbb{S},\preccurlyeq,\partial)$.
@@
Indeed, the pseudo-gap of $(\mathbb{S},\preccurlyeq,\partial)$ is either the maximum of $\Psi$ if this maximum exists, or the unique $\gamma\in\Gamma$ with $\Psi<\gamma<\{\gamma^{\prime}:\gamma\in\Gamma\wedge\gamma>0\}$ if $\Psi$ has no maximum ([4, Theorem 9.2.1 and Corollary 9.2.4]) and such an element exists.
@@
We say that asymptotic integration is regular on $\mathbb{S}$ if for all $d\in D$, there is a $d_{1}\in D$ with $d_{1}\geqslant d$ such that for each $f\in\mathbb{S}_{d}$ whose valuation is not the pseudo-gap of $(\mathbb{S},\preccurlyeq,\partial)$, there is a term $\tau\in C\mathfrak{M}$ with $\tau^{\prime}\sim f$ and $\tau^{\prime}\in\mathbb{S}_{d}$. We say that $\mathbb{S}$ has regular asymptotic integration if it is closed under asymptotic integration and asymptotic integration is regular on $\mathbb{S}$.
@@ 2 Conjugacy of derivations
The solutions are of the form $gA$ where $A^{\prime}\sim\frac{-\delta}{g^{2}}$. Thus $f$ and $g$ are asymptotically conjugate if and only if $-\delta/g^{2}$ has an asymptotic integral $A$ in $\mathbb{S}$ such that $gA$ lies in $\operatorname{Cont}(\partial)$. We conclude with Lemma 1.5. ∎
@@
[…]
@@
[…]
@@
By Theorem 2.1, there is an $h\in\operatorname{Cont}(\partial)$ with $(-h)\ast g\ast h=f$. Set $\sigma:=\exp(h\partial)$. Then $\sigma\circ\exp(g\partial)\circ\sigma^{\operatorname{inv}}=\exp(f\partial)$. Now the conjugation by $\sigma$ is a strongly linear algebra automorphism of $\operatorname{Der}^{\mathcal{S}}_{\prec}(\mathbb{S})$ (see [9, Proposition 1.28]), so $f\partial=\log(\sigma\circ\exp(g\partial)\circ\sigma^{\operatorname{inv}})=\sigma\circ\log(\exp(g\partial))\circ\sigma^{\operatorname{inv}}=\sigma\circ(g\partial)\circ\sigma^{\operatorname{inv}}$. ∎

### v4

@@ Abstract
We study conjugacy of formal derivations on fields of generalised power series in characteristic 0 and dimension $1$. Casting the problem of Poincaré resonance in terms of asymptotic differential algebra, we give conditions for conjugacy of parabolic flat $\log$-$\exp$ transseries, flat grid-based transseries, logarithmic transseries, power series with exponents and coefficients in an ordered field, and formal Puiseux series.
@@ Introduction
In this paper, we focus on the conjugacy problem for parabolic series in differential valued fields [34]. These include for instance formal Laurent or Puiseux series, $\log$-$\exp$ transseries or grid-based transseries [24], logarithmic transseries [21], or $\omega$-series [5], and complexifications thereof. This choice is not fortuitous but motivated by the connections between the setting of (ordered) asymptotic differential algebra and that of (ordered) valued groups (see [6, Remark 7.27]). One of the difficulties of solving conjugacy equations for formal series endowed with a composition law $\circ$ and a derivation $\partial$ is that this requires a good understanding of the interaction of the composition law with the valuation. On this path, one is confronted with intricate and computationally heavy problems involving Taylor expansions of arbitrarily high orders, monotonicity of the composition law, and mean value inequalities. It is not the least of hindrances that such properties of $\circ$ and $\partial$ may not have been established for the given algebra of formal series.
We circumvent these issues by leveraging the Lie-type correspondence, given by a formal exponential map $\exp$, between near-identity substitutions $f\mapsto f\circ(x+\delta)$ and contracting derivations $f\mapsto g\partial(f)$ on algebras of formal series. We showed [9] that a fraction of the theory of Lie groups applies to such algebras. We focus on contracting derivations on fields of generalised series, i.e. derivations which piecewise increase the valuation of non-trivial elements. On the side of vector fields, contractiveness can be related to nilpotency (see [28]).
@@
[Theorem 2.1] Suppose that $\mathbb{S}$ has regular asymptotic integration. Let $f,g\in\operatorname{Cont}(\partial)\setminus\{0\}$. Then $f$ and $g$ are conjugate in $\operatorname{Cont}(\partial)$ if and only if v(fg)>v(g)v(f-g)>v(g) and $v(f-g)>v\left(\left(h/gh^{\prime}\right)^{\prime}\right)$ for all $h\in\mathbb{S}^{\times}$ with $v(h)>0$.
@@
[Theorem 2.2] Suppose that $\mathbb{S}$ has regular asymptotic integration. Let $f,g\in\operatorname{Cont}(\partial)\setminus\{0\}$ such that v(fg)>v(f)v(f-g)>v(f) and and $v(f-g)>v\left(\left(h/gh^{\prime}\right)^{\prime}\right)$ for all $h\in\mathbb{S}^{\times}$ with $v(h)>0$. Then the derivations $f\partial:=h\mapsto f\partial(h)$ and $g\partial:=h\mapsto g\partial(h)$ are conjugate over $\operatorname{Aut}(\mathbb{S})$, i.e. there is a $\sigma=\exp(h\partial)\in\operatorname{Aut}(\mathbb{S})$ such that $\sigma\circ(g\partial)\circ\sigma^{\operatorname{inv}}=f\partial$.
In certain cases, the group of automorphisms $\exp(\operatorname{Cont}(\partial))$ is isomorphic to a well-identified group of series under composition. Let $\mathbb{S}$ be the field of transseries whose transmonomials $\mathfrak{m}$ satisfy $v(\mathfrak{m}^{\prime}/\mathfrak{m})+v(x)\geqslant 0$. We identify (Proposition 3.1) the group $\exp(\operatorname{Cont}(\partial))$ for all direct limits of subsystems (Definition 3.1) of the direct system which defines $\mathbb{S}$. Combining this with Theorem 1, we obtain:
@@ 1 Groups of contracting derivations
Throughout the paper, we fix a field $C$ of characteristic 0, a non-empty directed set $(D,\leqslant)$ and a directed system $\mathcal{S}=(\Gamma_{d})_{d\in D}$, for the inclusion, of non-trivial ordered Abelian groups. We write $\Gamma$ for the direct limit of $\Gamma_{d}$.
@@
Lastly, we write $1$-$\operatorname{Aut}^{\mathcal{S}}(\mathbb{S})$ for the group, under composition, of regular algebra automorphisms $\sigma$ of $\mathbb{S}$ such that $\sigma(f)\sim f$ for all $f\in\mathbb{S}$. Since each $C\operatorname{Id}_{\mathbb{S}_{d_{0}}}+\operatorname{Lin}^{+}_{\prec}(\mathbb{S}_{d_{0}})$ has evaluations in the sense of [9, Definition 2.3], and in view of the definition of regularity, so has $C\operatorname{Id}_{\mathbb{S}}+\operatorname{Lin}^{\mathcal{S}}_{\prec}(\mathbb{S})$. So [9, Theorem 2.14] applies and yields:
@@
$\exp:(\operatorname{Der}^{\mathcal{S}}_{\prec}(\mathbb{S}),\ast)$ $\longrightarrow$ $(1\text{-}\operatorname{Aut}^{\mathcal{S}}(\mathbb{S}),\circ)$
$\partial$ $\longmapsto$ $\operatorname{Id}_{\mathbb{S}}+\partial+\frac{1}{2}\partial\circ\partial+\frac{1}{6}\partial\circ\partial\circ\partial+\cdots.$
Hidden terms in (1) are $\mathbb{Q}$-linear combinations of iterated Lie brackets of lengths >3>3.
@@ 1.1 Contractive hull of a derivation
The inverse of an $f\in\operatorname{Cont}(\partial)$ for $\ast$ is simply f-f. We also have the following consequences of [6, Lemmas 7.14 and 7.15 and Remark 7.19]. We give the proofs here for completion.
@@
[…]
@@
We have $\llbracket f,g\rrbracket=\llbracket f-g,g\rrbracket$, so this follows from Lemma 1.1. ∎
In view of (2), Lemmas 1.1 and 1.2 and Remark 1.1, we obtain:
Corollary 1.1
@@
If $D=\{\bullet\}$, $\Gamma_{\bullet}=\Gamma=(\mathbb{Z},+,0,<)$, and $\partial=\frac{\mathrm{d}}{\mathrm{d}t}$ is the derivation with respect to $t$ on $C(\!(t)\!)=C(\!(\mathbb{Z})\!)$, then $\operatorname{Cont}(\partial)=\left\{f\in C(\!(\mathbb{Z})\!):f\prec t^{2}\right\}$.
@@ 1.2 Integration and asymptotic integration
[…]
@@
[…]
@@
A crucial property of $(\mathbb{S},\preccurlyeq,\partial)$ is that [4, Theorem 9.2.1] there is at most one $\gamma\in\Gamma$ such that $\gamma\not\in(\Gamma\setminus\{0\})^{\prime}$. If such an element $\gamma$ exist, then we call it the pseudo-gap of $(\mathbb{S},\preccurlyeq,\partial)$. If no such $\gamma$ exists, then $\mathbb{S}$ is closed under asymptotic integration. Indeed, since $\operatorname{Ker}(\partial)=C$, for any $f\in\mathbb{S}\setminus\{0\}$ and any $g$ with $v(g^{\prime})=v(f)$, the element $cg$ is an asymptotic integral of $f$ where $c$ is the leading coefficient of $f/g^{\prime}$.
@@
Indeed, the pseudo-gap of $(\mathbb{S},\preccurlyeq,\partial)$ is either the maximum of $\Psi$ if this maximum exists, or the unique $\gamma\in\Gamma$ with $\Psi<\gamma<\{\gamma^{\prime}:\gamma\in\Gamma\wedge\gamma>0\}$ if $\Psi$ has no maximum ([4, Theorem 9.2.1 and Corollary 9.2.4]) and such an element exists. We will require the following technical lemma:
Lemma 1.5
Let $d_{0},d\in D$ with $d_{0}\leqslant d$ and let $\gamma\in\Gamma_{d}$ such that $\gamma^{\prime}$ is the pseudo-gap of $(\mathbb{S}_{d_{0}},\preccurlyeq,\partial)$. Then we have $\mathbb{N}\gamma+\Gamma_{d_{0}}\subseteq(\mathbb{N}\gamma+\Gamma_{d_{0}})^{\prime}$.
Proof.
Let $\alpha\in\Gamma_{d_{0}}$ and $n\in\mathbb{N}$. If $\alpha\notin(\Gamma_{d_{0}})^{\prime}$, then $\alpha=\gamma^{\prime}$. Since $n+1\neq 0$, we have $n\gamma+\alpha=((n+1)\gamma)^{\prime}\in(\mathbb{N}\gamma)^{\prime}$. Suppose now that $\alpha=\beta^{\prime}$ for a $\beta\in\Gamma_{d_{0}}$. If $n=0$ then $n\gamma+\alpha=\beta^{\prime}\in\Gamma_{d_{0}}^{\prime}$. We now assume that $n>0$. We have $\gamma^{\prime}\geqslant(\Gamma_{d_{0}}^{\neq})^{{\dagger}}$ and $\gamma<0$, so $\gamma^{{\dagger}}=\gamma^{\prime}-\gamma>(\Gamma_{d_{0}}^{\neq})^{{\dagger}}$. In particular $\gamma^{{\dagger}}>\beta^{{\dagger}}$, whence $(n\gamma)^{{\dagger}}=\gamma^{{\dagger}}>\beta^{{\dagger}}$. Thus $(n\gamma)^{\prime}+\beta>n\gamma+\beta^{\prime}=n\gamma+\alpha$, which means in view of the Leibniz rule that $(n\gamma+\beta)^{\prime}=n\gamma+\alpha$, so $n\gamma+\alpha\in\Gamma_{d}^{\prime}$.∎
@@
Remark 1.3
If $f,g\in\mathbb{S}$ and $(f_{i})_{i\in I}$ is a summable family in $\mathbb{S}$, then $\operatorname{supp}fg$ is contained in $\operatorname{supp}f+\operatorname{supp}g$, and $\operatorname{supp}\sum\limits_{i\in I}f_{i}$ is contained in $\bigcup\limits_{i\in I}\operatorname{supp}f_{i}$ (see [23, Section 3]). It follows that for $f,g\in\operatorname{Cont}(\partial)$ and for any submonoid $M\subseteq\Gamma$ with $\operatorname{supp}f\cup\operatorname{supp}g\subseteq M$ and $\partial(C(\!(M)\!))\subseteq C(\!(M)\!)$, the support of $f\ast g$ is contained in $M$.
We now define our notion of compatibility between asymptotic integration and the directed system. We will see that it is satisfied in most natural fields of generalised power series, such as transseries.
@@
We say that asymptotic integration is regular on $\mathbb{S}$ if for all $d_{0}\in D$, there is a $d\in D$ with $d\geqslant d_{0}$ such that for each $f\in\mathbb{S}_{d_{0}}$ whose valuation is not the pseudo-gap of $(\mathbb{S},\preccurlyeq,\partial)$, there is a term $\tau\in C\mathfrak{M}_{d}$ with $\tau^{\prime}\sim f$ and $\tau^{\prime}\in\mathbb{S}_{d_{0}}$. We then say that such a $d$ is associated to $d_{0}$. We say that $\mathbb{S}$ has regular asymptotic integration if it is closed under asymptotic integration and asymptotic integration is regular on $\mathbb{S}$.
@@ 2 Conjugacy of derivations
The solutions are of the form $gA$ where $A^{\prime}\sim\frac{-\delta}{g^{2}}$. Thus $f$ and $g$ are asymptotically conjugate if and only if $\delta/g^{2}$ has an asymptotic integral $A$ in $\mathbb{S}$ such that $gA$ lies in $\operatorname{Cont}(\partial)$. We conclude with Lemma 1.6. ∎
@@
Let $d_{0}\in D$ such that $f,g\in\mathbb{S}_{d_{0}}$, and let $d\geqslant d_{0}$ be as in the definition of regular asymptotic integration, with respect to $d_{0}$. If there is a $\mathfrak{u}\in\mathfrak{M}_{d}$ such that $v(\mathfrak{u}^{\prime})$ is the pseudo-gap of $(\mathbb{S}_{d_{0}},\preccurlyeq,\partial)$, and $\mathfrak{u}^{\prime}\in\mathbb{S}_{d_{0}}$, then we set $\mathfrak{m}:=\mathfrak{u}$. If not, we set $\mathfrak{m}:=1$. By induction on an ordinal $\alpha$, we construct a strictly $\prec$-decreasing sequence $(\tau_{\gamma})_{\gamma<\alpha}$ of terms in $\operatorname{Cont}(\partial)\cap(C(\!(\Gamma_{d_{0}}+\mathbb{N}v(\mathfrak{m}))\!))$ such that defining $\varphi_{\eta},\eta\leqslant\alpha$ as above, the sequence $(\varphi_{\eta}\ast g\ast(-\varphi_{\eta})-f)_{\eta\leqslant\alpha}$ is strictly $\prec$-decreasing and that for all $\gamma<\alpha$ and all $\varepsilon\prec\tau_{\gamma}$, we have $(\varphi_{\gamma+1}+\varepsilon)\ast g\ast(-\varphi_{\gamma+1}-\varepsilon)-f\prec\varphi_{\gamma}\ast g\ast(-\varphi_{\gamma})-f$. The $\varepsilon$ here is a stand-in for $\varphi_{\alpha}-\varphi_{\eta}$ for various values of $\eta<\alpha$.
Let $\alpha$ such that for all $\eta<\alpha$, the sequence $(\tau_{\gamma})_{\gamma<\eta}$ is defined and satisfies the conditions. We first make a few observations. Let $\gamma\leqslant\alpha$ and write $g_{\gamma}:=\varphi_{\gamma}\ast g\ast(-\varphi_{\gamma})$. We have $g\sim g_{\gamma}$ by (3), so $\frac{f-g_{\gamma}}{g^{2}}$ and $\frac{f-g_{\gamma}}{g_{\gamma}^{2}}$ have the same symptotic integrals. Recall that $(f-g_{\rho})_{\rho\leqslant\alpha}$ is strictly $\prec$-decreasing. So given $\rho_{1}\leqslant\rho_{2}\leqslant\alpha$, if $\frac{f-g_{\rho_{i}}}{g^{2}}$ has an symptotic integral $A_{i},i\in\{1,2\}$, then we have $v(A_{1})+v(g_{\rho_{1}})+\Psi>0\Longrightarrow v(A_{2})+v(g_{\rho_{2}})+\Psi>0$. Thus, by Lemma 2.1, the series $f$ and $g_{\gamma}$ are asymptotically conjugate if and only if $v(\frac{f-g_{\gamma}}{g^{2}})$ is not the pseudo-gap of $(\mathbb{S},\preccurlyeq,\partial)$.
Let us now give the inductive proof. If $\alpha$ is a limit, including the case when $\alpha=0$, then we need only check that we have $(\varphi_{\gamma+1}+\varepsilon)\ast g\ast(-\varphi_{\gamma+1}-\varepsilon)\prec g_{\alpha}-f$ for all $\gamma<\alpha$ and $\varepsilon\prec\tau_{\gamma}$. But this follows by induction, noting that $\varphi_{\eta}-\varphi_{\gamma+1}\prec\tau_{\gamma}$ for all $\gamma<\eta<\alpha$. Suppose that $\alpha=\eta+1$ is a successor ordinal. If $g_{\eta}=f$, then, setting $\lambda:=\alpha$, we are done. Suppose that $g_{\eta}\neq f$. If $\mu:=v\left(\frac{f-g_{\eta}}{g^{2}}\right)$ is the pseudo-gap of $(\mathbb{S},\preccurlyeq,\partial)$, then we set $\lambda:=\alpha$ and we are done.
Suppose now that $\mu$ is not the pseudo-gap of $(\mathbb{S},\preccurlyeq,\partial)$. Since $\mathfrak{m}^{\prime}\in\mathbb{S}_{d_{0}}$, Remark 1.3 applies, and the set $\operatorname{Cont}(\partial)\cap(C(\!(\Gamma_{d_{0}}+\mathbb{N}v(\mathfrak{m}))\!))$ is closed under $\ast$. We see with Lemma 1.5 that $v(\frac{f-g_{\eta}}{g^{2}})$ is not a pseudo-gap in $(\mathbb{S}_{d},\preccurlyeq,\partial)$, so by regularity of asymptotic integration, there is a term $\tau$ in $\mathbb{S}_{d}$ with $\tau^{\prime}\in\mathbb{S}_{d}$ and $\tau^{\prime}\sim-\frac{f-g_{\eta}}{g^{2}}$. Note that $\tau$ is unique, and in view of Lemma 1.5, we have $\tau\in C^{\times}\mathfrak{M}_{d_{0}}\mathfrak{m}^{\mathbb{N}}$. So we may extend the sequence by setting $\tau_{\eta}:=\operatorname{lead}(g)\tau$, which by Lemma 2.1 lies in $\operatorname{Cont}(\partial)$. We have $\tau^{\prime}\prec(\frac{\tau_{\gamma}}{\operatorname{lead}(g)})^{\prime}$ for all $\gamma<\eta$ since $(f-g_{\gamma})_{\gamma<\alpha}$ is strictly $\prec$-decreasing. We deduce with Lemma 1.4 that $\tau\prec\frac{\tau_{\gamma}}{\operatorname{lead}(g)}$, whence $\tau_{\eta}\prec\tau_{\gamma}$ for all $\gamma<\eta$.
Thus $\varphi_{\alpha}$ is well-defined, and it remains to show that the conditions related to that sequence are satisfied. By Lemma 2.1 with $(g_{\eta},f)$ in the role of $(f,g)$, we have
@@
for all $\varepsilon\in\mathbb{S}$ with $\varepsilon\prec\tau_{\eta}$, as claimed. The sequence $(\tau_{\gamma})_{\gamma<\alpha}$ is strictly $\prec$-decreasing, hence injective. Since $\mathbb{S}_{d}$ is a set, there is an ordinal $\lambda$ at which the process stops, i.e. one of the cases of the lemma occurs. ∎
Note that the recularity of asymptotic integration plays a crucial role in the previous proof in allowing us to remain in $\mathbb{S}_{d}$ within the induction, so that $\varphi_{\lambda}$ be well-defined.
@@
By Theorem 2.1, there is an $h\in\operatorname{Cont}(\partial)$ with $(-h)\ast g\ast h=f$. Set $\sigma:=\exp(h\partial)$. Then $\sigma\circ\exp(g\partial)\circ\sigma^{\operatorname{inv}}=\exp(f\partial)$. Now the conjugation by $\sigma$ is a strongly linear algebra automorphism of $\operatorname{Der}^{\mathcal{S}}_{\prec}(\mathbb{S})$ (see [9, Proposition 1.28]), while $\log$ is given as a power series of its argument with constant coefficients. So $f\partial=\log(\sigma\circ\exp(g\partial)\circ\sigma^{\operatorname{inv}})=\sigma\circ\log(\exp(g\partial))\circ\sigma^{\operatorname{inv}}=\sigma\circ(g\partial)\circ\sigma^{\operatorname{inv}}$. ∎